# Zeke Xie HKUST(GZ) Guangzhou, China

### v1

@@ I Introduction
Standard training from ambiguous data with such biased hard labels induces overconfident predictions [9, 10, 11, 6, 12, 13]. A related line of research on biased labels is learning with noisy labels [14], which generally studies unambiguous instances with randomly flipped noisy labels. Obviously, this classical label noise setting cannot reflect label bias from data ambiguity and human uncertainty. Therefore, a formal formulation and a novel method for learning from ambiguous data with hard labels are very much needed to address real-world challenges.
@@ II Quantized Label Learning
We call the process of mapping the label distribution to a single hard label label quantization and propose the label quantization hypothesis: the human-annotated label of an instance is sampled from its associated ground-truth soft-label distribution.
Following this, we formally define a quantized label in Definition 1.
@@ 14       Update the model parameters by a optimizer;
The choice of the loss function $l$. As for the choice of loss function $l$, a number of statistical distance/similarity measures may serve as alternatives, such as Cross Entropy (CE), Mean Absolute Errors (MAE), Mean Squared Errors (MSE), and Kullback–Leibler (KL) divergence. Different loss functions enjoy different advantages and disadvantages in the PU framework [17].
In this work, we used a stochastic variant of Jensen-Shannon (JS) divergence, named Stochastic JS (SJS) divergence, as $D_{\rm SJS}(P\|Q)=\alpha D_{\mathrm{KL}}(P\|M)+(1-\alpha)D_{\mathrm{KL}}(Q\|M),$ where $M=\alpha P+(1-\alpha)Q$. We note that $\alpha\in(0,0.5]$ obeys a Beta distribution, $\mathrm{Beta}(0.5,0.5)$, and is randomly sampled per iteration. SJS is an interpolated generalization of KL divergence (with $\alpha\to 0$ ) and JS divergence (with $\alpha=\frac{1}{2}$). Following [23], we recommend to use the scaled form $\hat{D}_{\rm SJS}(P\|Q)=-1/[(1-\alpha)\log(1-\alpha)]D_{\rm SJS}(P\|Q)$ in practice to make the loss magnitude relatively robust to $\alpha$ during training.
@@ IV Ambiguous Data Benchmark
By linear interpolation of instances, Mixup generates intrinsic ambiguous data which confuses annotators, especially for localization. In this case, the Mixup operation is expressed as $\operatorname{Mixup}({\bm{x}}_{1},\dots,{\bm{x}}_{m};\bm{\lambda})=\sum_{i=1}^{m}\lambda_{i}{\bm{x}}_{i}$. In the Mixup-based strategy, we let $r\bm{\lambda}$ obey a $m$-category multinominal distribution, namely, $r\bm{\lambda}\sim{\rm MultiNom}(1/m,\dots,1/m;r)$. We note that $r$ is the hyperparameter that controls the number of trials of the multinominal distribution and $\lambda_{i}$ indicates the frequency that the outcome $i$ occurs in $r$ trials.
Similarly, we generalize the original CutMix as a novel PatchMix, expressed as $\operatorname{PatchMix}({\bm{x}}_{1},\dots,{\bm{x}}_{m};\bm{\lambda})=\sum_{i=1}^{m}{\mathbf{M}}_{i}\odot{\bm{x}}_{i}$, where each binary mask ${\mathbf{M}}_{i}$ for ${\bm{x}}_{i}$ indicates which patches to drop out and fill in from a candidate input with the constrain $\sum_{i=1}^{m}{\mathbf{M}}_{i}=\bm{1}$, where $\bm{1}$ is an all-ones matrix.
@@ V Empirical Analysis and Discussion
In this section, we conduct extensive experiments to evaluate the effectiveness of the proposed CPU risk for QLL using highly ambiguous data.
Experimental Settings. We use CIFAR-10 and CIFAR-100 as the standard dataset for preparing ambiguous datasets with quantized labels, namely, CIFAR-10Q and CIFAR-100Q, following a Mixup-based method with $m=2$ and a PatchMix-based method with $m=4$ proposed in Section IV. We train ResNet-18 [26] via SGD on the ambiguous training dataset. We use the original test sets of CIFAR datasets for evaluation.
In practice, we may choose the class-prior hyperparameters $\pi_{{\textnormal{p}}}^{(1)}$ and $\pi_{{\textnormal{p}}}^{(2)}$ simply as $\pi_{{\textnormal{p}}}^{(1)}=0.1$ and $\pi_{{\textnormal{p}}}^{(2)}=\frac{m}{c}$ without fine-tuning costs in our experiments, as the choice consistently shows good performance in our experiments.
1. Various loss functions/risk estimators as the baselines. The proposed QLL problem setting does not have baseline methods. Given that quantized labels provide erroneous supervision for learning models, we thereby mainly compare CPU with multiple label-noise-robust baselines, including CE, Bootstrap (BS) [27], Generalized Cross Entropy (GCE) [28], Symmetric Cross Entropy (SCE) [29], Determinant-based Mutual Information (DMI) [30], and JS divergence [23]. CE is the most common training objective/risk in standard training. Bootstrap is a famous label-noise-robust method for conventional synthetic label noise. GCE is a recent label-noise-robust generalization of CE and Mean Absolute Error. SCE is the symmetric variant of CE which is also robust to label noise.
These methods mainly used specifically designed loss functions or risk estimators. Thus, they are directly comparable with the proposed CPU.

### v2

# Zeke Xie* $\dagger$ HKUST(GZ) Guangzhou, China

@@ I Introduction
Standard training from ambiguous data with such biased hard labels induces overconfident predictions [9, 10, 11, 12, 13, 14, 15, 16]. A related line of research on biased labels is learning with noisy labels [17], which generally studies unambiguous instances with randomly flipped noisy labels. Obviously, this classical label noise setting cannot reflect label bias from data ambiguity and human uncertainty. Therefore, a formal formulation and a novel method for learning from ambiguous data with hard labels are very much needed to address real-world challenges.
@@ II Quantized Label Learning
We call the process of mapping the label distribution to a single hard label label quantization and propose the label quantization hypothesis: the human-annotated label of an instance is sampled from its associated ground-truth soft-label distribution. Following this, we formally define a quantized label in Definition 1.
@@ 14       Update the model parameters by a optimizer;
The choice of the loss function $l$. As for the choice of loss function $l$, a number of statistical distance/similarity measures may serve as alternatives, such as Cross Entropy (CE), Mean Absolute Errors (MAE), Mean Squared Errors (MSE), and Kullback–Leibler (KL) divergence. Different loss functions enjoy different advantages and disadvantages in the PU framework [20]. In this work, we used a stochastic variant of Jensen-Shannon (JS) divergence, named Stochastic JS (SJS) divergence, as $D_{\rm SJS}(P\|Q)=\alpha D_{\mathrm{KL}}(P\|M)+(1-\alpha)D_{\mathrm{KL}}(Q\|M),$ where $M=\alpha P+(1-\alpha)Q$. We note that $\alpha\in(0,0.5]$ obeys a Beta distribution, $\mathrm{Beta}(0.5,0.5)$, and is randomly sampled per iteration. SJS is an interpolated generalization of KL divergence (with $\alpha\to 0$ ) and JS divergence (with $\alpha=\frac{1}{2}$). Following [26], we recommend to use the scaled form $\hat{D}_{\rm SJS}(P\|Q)=-1/[(1-\alpha)\log(1-\alpha)]D_{\rm SJS}(P\|Q)$ in practice to make the loss magnitude relatively robust to $\alpha$ during training.
@@ IV Ambiguous Data Benchmark
By linear interpolation of instances, Mixup generates intrinsic ambiguous data which confuses annotators, especially for localization. In this case, the Mixup operation is expressed as $\operatorname{Mixup}({\bm{x}}_{1},\dots,{\bm{x}}_{m};\bm{\lambda})=\sum_{i=1}^{m}\lambda_{i}{\bm{x}}_{i}$. In the Mixup-based strategy, we let $r\bm{\lambda}$ obey a $m$-category multinominal distribution, namely, $r\bm{\lambda}\sim{\rm MultiNom}(1/m,\dots,1/m;r)$. We note that $r$ is the hyperparameter that controls the number of trials of the multinominal distribution and $\lambda_{i}$ indicates the frequency that the outcome $i$ occurs in $r$ trials. Similarly, we generalize the original CutMix as a novel PatchMix, expressed as $\operatorname{PatchMix}({\bm{x}}_{1},\dots,{\bm{x}}_{m};\bm{\lambda})=\sum_{i=1}^{m}{\mathbf{M}}_{i}\odot{\bm{x}}_{i}$, where each binary mask ${\mathbf{M}}_{i}$ for ${\bm{x}}_{i}$ indicates which patches to drop out and fill in from a candidate input with the constrain $\sum_{i=1}^{m}{\mathbf{M}}_{i}=\bm{1}$, where $\bm{1}$ is an all-ones matrix.
@@ V Empirical Analysis and Discussion
In this section, we conduct extensive experiments to evaluate the effectiveness of the proposed CPU risk for QLL.
Experimental Settings. We use CIFAR-10/100 as the standard dataset for preparing ambiguous datasets with quantized labels, namely, CIFAR-10Q and CIFAR-100Q, following a Mixup-based method with $m=2$ and a PatchMix-based method with $m=4$ proposed in Section IV. We train ResNet-18 [29] via SGD on the ambiguous training dataset. We use the original test sets of CIFAR datasets for evaluation. In practice, we may choose the class-prior hyperparameters simply as $\pi_{{\textnormal{p}}}^{(1)}=0.1$ and $\pi_{{\textnormal{p}}}^{(2)}=\frac{m}{c}$ without fine-tuning in our work.
1. Various loss functions/risk estimators as the baselines. The proposed QLL problem setting does not have baseline methods. Given that quantized labels provide erroneous supervision for learning models, we thereby mainly compare CPU with multiple label-noise-robust baselines, including CE, Bootstrap (BS) [30], Generalized Cross Entropy (GCE) [31], Symmetric Cross Entropy (SCE) [32], Determinant-based Mutual Information (DMI) [33], and JS divergence [26]. CE is the most common training objective/risk in standard training. Bootstrap is a famous label-noise-robust method for conventional synthetic label noise. GCE is a recent label-noise-robust generalization of CE and Mean Absolute Error. SCE is the symmetric variant of CE which is also robust to label noise. These methods mainly used designed loss functions or risk estimators, which are directly comparable to CPU.
@@ VI Conclusion
Acknowledgment
This work was supported by Science and Technology Bureau of Nansha District under Key Field Science and Technology Plan Program No. 2024ZD002 and Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No.2023B1212010007).